# Unlearning via Sparse Representations

### v1

@@ Abstract
Machine unlearning, which involves erasing knowledge about a forget set from a trained model, can prove to be costly and infeasible by existing techniques. We propose a nearly compute-free zero-shot unlearning technique based on a discrete representational bottleneck. We show that the proposed technique efficiently unlearns the forget set and incurs negligible damage to the model’s performance on the rest of the data set. We evaluate the proposed technique on the problem of class unlearning using three datasets: CIFAR-10, CIFAR-100, and LACUNA-100. We compare the proposed technique to SCRUB, a state-of-the-art approach which uses knowledge distillation for unlearning. Across all three datasets, the proposed technique performs as well as, if not better than SCRUB while incurring almost no computational cost.
@@ 1 Introduction
Machine Unlearning (Cao & Yang, 2015; Nguyen et al., 2022; Zhang et al., 2023; Xu et al., 2023; Kurmanji et al., 2023) may be defined as the problem of removing the influence of a subset of the data on which a model has been trained. Unlearning can be an essential component in addressing several problems encountered in deploying deep-learning-based solutions in real life. Neural networks such as Large Language Models (LLMs), which have been trained on massive amounts of commonly available data, can exhibit harmful behaviors in the form of generating misinformation, demonstrating harmful biases, or having other undesirable characteristics. A major culprit behind these behaviors is the presence of biased or corrupted instances in the training data of these models. To ensure safe model deployment, it is necessary to remove these instances. Another reason to remove instances and make a model behave as if it had not been trained on certain data is concerns about data privacy and the right of end users to expunge their data (Mantelero, 2013). For example, an individual might want their data removed from a face recognition system that was trained on their faces such that it is no longer able to identify them. All the above problems can be addressed by unlearning a specific subset of the training data, i.e., the subset of data giving rise to the harmful behavior of the model in the former cases and an individual’s private data in the latter cases. Apart from these concerns, unlearning can also be used for other purposes such as removing outdated data from a model to free the network capacity for more recent or relevant data. With increasing concerns about AI safety and the increasing ubiquity of deep learning models in real-world applications, the problem of unlearning is becoming critical.
The main challenge in unlearning is maintaining the performance of the model on the data that needs to be retained, called the retain set, while unlearning the forget set. The naive way to ensure that a model has no information about the forget set is to train from scratch on the retain set. Unlearning techniques aim to achieve the same end but at a much lower computational cost compared to full retraining. Unlearning in a pretrained network is difficult, especially in densely connected neural networks, since the value of one parameter may affect the output for all the input examples given to the neural network. A possible solution is to fine tune the model we wish to unlearn only on the retain set. While this would ensure that the performance of the model on the retain set is maintained, it has been shown to be ineffective in practice in unlearning the forget set (Golatkar et al., 2020a). Other more effective solutions include retraining the model on the training data with a negative gradient for the forget set (Golatkar et al., 2020a; Kurmanji et al., 2023), or using knowledge-distillation-based training objectives to capture information about the retain set while filtering out information about the forget set (Kurmanji et al., 2023; Chundawat et al., 2023). However, all of these approaches require some form of substantial additional compute in order to facilitate unlearning. Moreover, some of the existing approaches also require access to the original training data to enable unlearning, which may not be possible in many practical applications, e.g., for a model in production which is being trained online on an incoming data stream. The use of large models is becoming more popular and prevalent with the development of general purpose transformer models. The requirement for additional compute can quickly become impractical in the context of these large models, especially in cases where a model is deployed and needs to be redeployed as quickly as possible after making the necessary changes.
@@
Figure 1: A summary of the proposed unlearning approach. Left: The structure of a key-value bottleneck. The encoder is frozen and pre-trained and $R_{1}$ is a random projection matrix. The values corresponding to the selected keys are retrieved to be used by the decoder. The gradient is backpropagated through the decoder into the values during training. The figure depicts the case with 1 codebook in the DKVB. However, in practice we use multiple codebooks. Center: Examples from the forget set are passed through the trained model and the key-value pairs selected during the forward pass are recorded. Right: The recorded key-value pairs are then masked from the bottleneck. As a result, the key selection is redirected to other non-masked keys, whose corresponding values are not useful for the example, and just lead to uninformed prediction.
@@
In this article, we argue that neural information bottlenecks can be means of highly efficient and specific unlearning. Neural information bottlenecks have emerged as useful components in neural network architectures, providing numerous benefits such as improving out-of-distribution (OOD) generalization capabilities and robustness to noisy data (Goyal et al., 2021; Jaegle et al., 2021; Liu et al., 2021; 2023), facilitating large scale unsupervised pre-training and generative modeling (Esser et al., 2021; Oord et al., 2017), and more recently, helping in continual learning (Träuble et al., 2023). We particularly focus on using the Discrete Key-Value Bottleneck (DKVB) proposed in Träuble et al. (2023). DKVB induces sparse representations in the form of key-value pairs which are trained in a localized and context-dependent manner. Since these representations are sparse, we hypothesize that it is possible to remove the information about a subset of the training data without damaging the information about the rest of the data—the primary desiderata for a useful unlearning method. Moreover, since the representations are discrete, this may be achieved zero-shot, i.e., without requiring any additional compute in the form of retraining or fine tuning, by directly intervening on individual representations.
@@
In this work, we investigate the above-mentioned idea of zero-shot unlearning in the Discrete Key-Value Bottleneck. Specifically, we focus on the problem of class unlearning in multi-class classification tasks, where the aim is to remove information about a specific output class, called the forget class, from a trained model. We use the term retain classes to refer to the classes other than the forget class that are present in the training data. More specifically, we wish to remove the influence of the forget class (or more generally speaking the forget set) on the model. We measure this influence using the performance of the model on held-out test datasets corresponding to the forget class and the retain classes. We propose two approaches for zero-shot unlearning on DKVB, one which requires access to training examples from the forget set—which we call Unlearning via Examples—and one which does not require access to training examples—called Unlearning via Activations. We show that the proposed methods achieve unlearning of the forget class while incurring negligible damage to the model’s performance on the retain classes. We compare the proposed methods to SCRUB (Kurmanji et al., 2023), a recent state-of-the-art approach that requires additional compute to unlearn, on three datasets: CIFAR-10, CIFAR-100, and LACUNA-100, and also further investigate the effects of retraining the zero-shot unlearned models on the retain set.
@@ 2 Related Work
The problem of unlearning has been studied in different forms for over two decades. Early works such as Tsai et al. (2014), Cauwenberghs & Poggio (2000) and Duan et al. (2007) study the problem of decremental learning in linear models, where a small number of samples need to be removed from a model. Ginart et al. (2019) considers unlearning as a problem of deleting individual data points from a model. They give a probabilistic definition, formalize the notion of efficient data deletion, and propose two deletion efficient learning algorithms. Guo et al. (2019) introduces certified removal - a theoretical guarantee of indistinguishability between a model from which data was removed and a model that never saw the data. Izzo et al. (2021) distinguishes between exact unlearning and approximate unlearning and proposes a compute-efficient approximate data deletion method, and a new metric for evaluating data deletion from these models. Golatkar et al. (2020a) and Kurmanji et al. (2023) cast unlearning into an information theoretic framework. Golatkar et al. (2020b) proposes Neural Tangent Kernel (NTK) (Jacot et al., 2018) theory-based approximation of the weights of the unlearned network. Multiple works also delve into the more philosophical, ethical, and legal aspects of unlearning and the “right to be forgotten” (Kwak et al., 2017; Villaronga et al., 2018). Chundawat et al. (2023); Tarun et al. (2023) learn error minimization and error maximization-based noise matrices which are used to finetune the trained model in order to do unlearning. Chundawat et al. (2023) further uses a generator that generates pseudo data points for unlearning in order to operate in a data-free regime.
@@
Most relevant to our work, Kurmanji et al. (2023) introduces SCRUB, an effective knowledge distillation-based unlearning method. SCRUB considers the original model as a teacher model and trains a student model to obey the teacher model on the retain set and disobey it on the forget set. This is done by computing the KL Divergence between the output distributions of the two models and training the student model to maximize it on the forget set (this is called a max-step) and minimize it on the retain set (this is called a min-step). The student model is simultaneously also optimized for minimizing the task loss on the retain set. The training consists of mstep max-steps. The max-steps and min-steps are performed in an interleaved fashion. Warnecke et al. (2021) focus on unlearning in deleting information at the features level rather than unlearning specific instances or classes. They do so by using an optimization objective that incentivizes a model to replace the information about a set of features with a perturbed set of features. The proposed approaches on the other hand focus on unlearning entire classes of data. Chen et al. (2023), similarly to us, focuses on class unlearning. Unlearning is done by destroying the decision boundary of the forget class. The authors propose two boundary shift methods termed as Boundary Shrink and Boundary Expanding.
@@
Jia et al. (2023) and Mehta et al. (2022) investigate machine unlearning in context of model sparsity in context of sparsity. Jia et al. (2023) leverages the Lottery Ticket Hypothesis Frankle & Carbin (2018) by using parameter pruning on a trained dense model to identify the token subnetwork. They observe that applying standard unlearning approaches to a sparsified networks is better as compared to doing unlearning directly on the dense network. Mehta et al. (2022) identify the Markovian Blanket of parameters corresponding to the examples to be unlearnt and updates those parameters. Thus, their approach can be seen as applying sparse updates to the network for unlearning. Both these approaches start with dense trained models and leverage sparsity for unlearning. Whereas the approaches proposed in this paper suggest using sparsity as an inductive bias in the model during the initial training along which combined with the architectural prior of the DKVB makes it suitable for unlearning involving minimal compute requirements. Jia et al. (2023) on the hand spasifies the model after it has been trained. Mehta et al. (2022) involves sparse updates to the model parameters as discussed previously. However, these sparse updates are utilized during unlearning as opposed to during training of the original model in the proposed approaches.
While most of the above-mentioned approaches improve upon the naive and intractable baseline of retraining on the retain set, in terms of computational efficiency, they still require some amount of computation to do unlearning. This additional compute requirement can quickly become infeasible where large models are involved. The proposed approach, on the other hand, requires negligible computation for unlearning. Any computation that may be required is in the form of running inference on the forget set.
@@ 3 Background and Notations
Unlearning: Let $\mathcal{D}_{train}=\{x_{i},y_{i}\}^{N}_{i=1}$ be a training dataset and $\mathcal{D}_{test}$ be the corresponding test dataset. In our experiments, we consider the setting of class unlearning, wherein we aim to unlearn a class $c$ from a model trained with a multiclass classification objective on $\mathcal{D}_{train}$. $c$ is called the forget class or the forget set. Given $c$, we obtain $\mathcal{D}_{train}^{forget}\subset\mathcal{D}_{train}$ such that $\mathcal{D}_{train}^{forget}=\{(x,y)\in\mathcal{D}_{train}|y=c\}$. The complement of $\mathcal{D}_{train}^{forget}$ is $\mathcal{D}_{train}^{retain}$, i.e., subset of $\mathcal{D}_{train}$ that we wish to retain. Thus $\mathcal{D}_{train}^{retain}\cup\mathcal{D}_{train}^{forget}=\mathcal{D}_{train}$. Similarly, from $\mathcal{D}_{test}$, we get $\mathcal{D}_{test}^{forget}=\{(x,y)\in\mathcal{D}_{test}|y=c\}$ and its complement $\mathcal{D}_{test}^{retain}$. We refer to $\mathcal{D}_{train}^{retain}$ as the retain set training data, $\mathcal{D}_{test}^{retain}$ as the retain set test data, $\mathcal{D}_{train}^{forget}$ as the forget set training data and $\mathcal{D}_{test}^{forget}$ as the forget set test data.
Discrete Key-Value Bottleneck: A discrete key-value bottleneck (DKVB) (Träuble et al., 2023) consists of a discrete set of coupled key-value codes. Specifically, the bottleneck contains $C$ codebooks with each codebook containing $M$ key-value pairs. Models with DKVB use a pre-trained and frozen encoder to encode the input into a continuous representation. This input representation is then projected into $C$ lower dimension heads and each head is quantized to the $top-k$ nearest keys in the corresponding codebook. The values corresponding to the selected keys are averaged, and used for the downstream task. The keys in the codebooks are frozen and initialized to cover the input data manifold whereas the values are learnable. The mapping between the keys and values is non-parametric and frozen. Thus, the gradient is not propagated between the values and keys during training of the model. Since the values are retrieved and updated sparsely, and all the components except the value codes and the decoder are frozen, DKVB stores information in the form input-dependent, sparse and localized representations (i.e., the value codes). These inductive biases allow the framework to exhibit improved generalization under distribution shifts during training, as shown empirically in Träuble et al. (2023). Figure 1 (Left) shows an overview of a model with a DKVB where $C=1$, $M=5$ and top-$k=1$.

### v2

# Low Compute Unlearning via Sparse Representations

@@ Abstract
Machine unlearning, which involves erasing knowledge about a forget set from a trained model, can prove to be costly and infeasible using existing techniques. We propose a low compute unlearning technique based on a discrete representational bottleneck. We show that the proposed technique efficiently unlearns the forget set and incurs negligible damage to the model’s performance on the rest of the data set. We evaluate the proposed technique on the problem of class unlearning using four datasets: CIFAR-10, CIFAR-100, LACUNA-100 and ImageNet-1k. We compare the proposed technique to SCRUB, a state-of-the-art approach which uses knowledge distillation for unlearning. Across all three datasets, the proposed technique performs as well as, if not better than SCRUB while incurring almost no computational cost.
@@ 1 Introduction
Machine Unlearning (Cao & Yang, 2015; Nguyen et al., 2022; Zhang et al., 2023; Xu et al., 2023; Kurmanji et al., 2023; Warnecke et al., 2021) may be defined as the problem of removing the influence of a subset of the data on which a model has been trained. Unlearning can be an essential component in addressing several problems encountered in deploying deep-learning approaches in practical scenarios. Neural networks such as Large Language Models (LLMs), trained on massive amounts of commonly available data, can exhibit harmful behaviors in the form of generating misinformation, demonstrating harmful biases, or other undesirable characteristics. A major culprit behind these behaviors is the presence of biased or corrupted instances in the training data of these models. To ensure safe model deployment, it is necessary to remove these instances. Another reason to remove instances and make a model behave as if it had not been trained on certain data is concerns about data privacy and the right of end users to expunge their data (Mantelero, 2013; Dang, 2021). For example, an individual might want their data removed from a face recognition system that was trained on their faces such that it is no longer able to identify them. Several regulations are being put in place in order to safeguard the ”right to be forgotten” (Pardau, 2018; Magdziarczyk, 2019). All the above problems can be addressed by unlearning ther specific subset of the training data which gives rise to the harmful behavior of the model in the former cases and an individual’s private data in the latter cases. Apart from these concerns, unlearning can also serve other purposes such as removing outdated data from a model to free the network capacity for more recent or relevant data. With increasing concerns about AI safety and the increasing ubiquity of deep learning models in real-world applications, the problem of unlearning is of critical importance.
@@
Figure 1: A summary of the proposed unlearning approach. Left: The structure of a key-value bottleneck. The encoder is frozen and pre-trained and $R_{1}$ is a random projection matrix. The values corresponding to the selected keys are retrieved to be used by the decoder. The gradient is backpropagated through the decoder into the values during training. The figure depicts the case with 1 codebook in the DKVB. However, in practice we use multiple codebooks. Center: Examples from the forget set are passed through the trained model and the key-value pairs selected during the forward pass are recorded. Right: The recorded key-value pairs are then masked from the bottleneck. As a result, the key selection is redirected to other keys, with non-informative corresponding values leading to other prediction.
The main challenge in unlearning is maintaining the performance of the model on the data that needs to be retained, called the retain set, while unlearning the forget set. The naive way to ensure that a model has no information about the forget set is to train from scratch on the retain set. Unlearning techniques aim to achieve the same goal but at a much lower computational cost compared to full retraining. Unlearning in a pretrained network is difficult, especially in densely connected neural networks, since the value of one parameter may affect the output for all the input examples given to the neural network. A possible solution is to fine tune the model we wish to unlearn only on the retain set. While this would ensure that the performance of the model on the retain set is maintained, it can be computationally infeasible in practice. Other more effective solutions include retraining the model on the training data with a negative gradient for the forget set (Golatkar et al., 2020a; Kurmanji et al., 2023), or using knowledge-distillation-based training objectives to capture information about the retain set while filtering out information about the forget set (Kurmanji et al., 2023; Chundawat et al., 2023). Nevertheless, all of these approaches require some form of substantial additional compute in order to facilitate unlearning. Moreover, some of the existing approaches additionally require access to the original training data to facilitate unlearning, which may not be possible in many practical applications, e.g., for a model in production which is being trained online on an incoming data stream. The use of large models is becoming more popular and prevalent with the advent of general purpose transformer models. The requirement for additional compute can quickly become impractical in the context of these large models, especially in cases where a model is deployed and needs to be redeployed as quickly as possible after making the necessary changes.
@@
In this article, we argue that specific kinds of discrete neural information bottlenecks are highly suited for very efficient and specific unlearning. Neural information bottlenecks have emerged as useful components in neural network architectures, providing numerous benefits such as improving out-of-distribution (OOD) generalization capabilities and robustness to noisy data (Goyal et al., 2021; Jaegle et al., 2021; Liu et al., 2021; 2023), facilitating large scale unsupervised pre-training and generative modeling (Esser et al., 2021; Oord et al., 2017), and more recently, helping in continual learning (Träuble et al., 2023). In particular, we build upon Discrete Key-Value Bottleneck (DKVB) proposed in Träuble et al. (2023). DKVB induces sparse representations in the form of key-value pairs which are trained in a localized and context-dependent manner. Since these representations are sparse, we hypothesize that it is possible to remove the information about a subset of the training data without damaging the information about the rest of the data—the primary desiderata for a useful unlearning method. Moreover, since the representations are discrete, this may be achieved without requiring any additional compute in the form of retraining or fine tuning, by directly intervening on individual representations.
@@
We investigate the above-mentioned idea of low compute unlearning in the Discrete Key-Value Bottleneck. Specifically, we focus on the problem of class unlearning in multi-class classification tasks, where the aim is to remove information about a specific class, called the forget class, from a trained model. We use the term retain classes to refer to the classes other than the forget class that are present in the training data. More specifically, we wish to remove the influence of the forget class on the model. We measure this influence using the performance of the model on held-out test datasets corresponding to the forget class and the retain classes.
We propose two approaches for compute efficient unlearning in DKVB - Unlearning via Examples and Unlearning via Activations. We show that the proposed methods achieve unlearning of the forget class while incurring negligible damage to the model’s performance on the retain classes. We compare the proposed methods to SCRUB (Kurmanji et al., 2023), a recent state-of-the-art approach that requires additional compute to unlearn, on four datasets: CIFAR-10, CIFAR-100, LACUNA-100 and ImageNet-1k. The novelty of our work lies in the largely under-explored idea of using a model architecture with inherent sparse representations for unlearning.
@@ 2 Related Work
The problem of unlearning has been studied in different forms for over two decades. Early works such as Tsai et al. (2014), Cauwenberghs & Poggio (2000) and Duan et al. (2007) study the problem of decremental learning in linear models, where a small number of samples need to be removed from a model. Ginart et al. (2019) considers unlearning as a problem of deleting individual data points from a model. They give a probabilistic definition, formalize the notion of efficient data deletion, and propose two deletion efficient learning algorithms. Guo et al. (2019) introduces certified removal - a theoretical guarantee of indistinguishability between a model from which data was removed and a model that never saw the data. Izzo et al. (2021) distinguishes between exact unlearning and approximate unlearning and proposes a compute-efficient approximate data deletion method, and a new metric for evaluating data deletion from these models. Golatkar et al. (2020a) and Kurmanji et al. (2023) cast unlearning into an information theoretic framework. Golatkar et al. (2020b) proposes Neural Tangent Kernel (NTK) (Jacot et al., 2018) theory-based approximation of the weights of the unlearned network. Multiple works also delve into the more philosophical, ethical, and legal aspects of unlearning and the “right to be forgotten” (Kwak et al., 2017; Villaronga et al., 2018). Chundawat et al. (2023) and Tarun et al. (2023) learn error minimization and error maximization-based noise matrices which are used to finetune the trained model in order to do unlearning. Chundawat et al. (2023) further uses a generator that generates pseudo data points for unlearning in order to operate in a data-free regime.
@@
Most relevant to our work, Kurmanji et al. (2023) introduces SCRUB, a knowledge distillation-based unlearning method. SCRUB considers the original model as a teacher model and trains a student model to obey the teacher model on the retain set and disobey it on the forget set. This is done by computing the KL Divergence between the output distributions of the two models and training the student model to maximize it on the forget set (called a max-step) and minimize it on the retain set (called a min-step). The student model is simultaneously also optimized for minimizing the task loss on the retain set. The training consists of mstep max-steps. The max-steps and min-steps are executed alternatively. Chen et al. (2023), similarly to us, focuses on class unlearning rather than unlearning specific instances in the data. Unlearning is done by destroying the decision boundary of the forget class. The authors propose two boundary shift methods termed as Boundary Shrink and Boundary Expanding.
@@
Jia et al. (2023) and Mehta et al. (2022) investigate unlearning in context of model sparsity. Jia et al. (2023) leverages the Lottery Ticket Hypothesis, Frankle & Carbin (2018) leverages using parameter pruning on a trained dense model to identify the token subnetwork. They observe that applying standard unlearning approaches to a sparsified networks is better as compared to doing unlearning directly on the dense network. Mehta et al. (2022) identifies the Markovian Blanket of parameters corresponding to the examples to be unlearnt and updates those parameters. Their approach can be seen as applying sparse unlearning updates to the network.
While most of the approaches discussed above improve upon the naive and intractable baseline of retraining on the retain set, all of them require a substantial amount of additional computation in the form of optimizing an objective function for unlearning. This additional compute requirement can quickly become infeasible whenever large models are involved. The proposed approach in this work, on the other hand, requires negligible computation for unlearning. Any computation that may be required is in the form of running inference on the forget set. We point out that sparsity is a critical dimension that determines the effectiveness of unlearning: extremely sparse representations make unlearning trivial, whereas fully distributed representations intertwine knowledge in a way that makes compute-efficient unlearning a serious challenge. Previous methods studying sparsity in the context of unlearning such as Jia et al. (2023) and Mehta et al. (2022) propose the use of pruning techniques to first sparsify the network. These approaches start with dense trained models and leverage sparsity for unlearning. In contrast, we propose using sparsity as an in-built inductive bias in the model during the initial training which makes the model suitable for unlearning involving minimal compute requirements. On the other hand, Jia et al. (2023) sparsify the model after it has been trained. Mehta et al. (2022) involves sparse updates to the model parameters as discussed previously. However, these sparse updates are utilized during unlearning as opposed to during training of the original model in the proposed approach. We identify a sweet spot on the continuum between local and distributed learning that allows for both, compute-efficient unlearning and simultaneously obtaining the same generalization performance.
Xu et al. (2023) have introduced a taxonomy that categorizes existing research on unlearning based on different approaches and aims. In this classification, our methods fall within the Model Pruning category by means of disabling specific (key, value) pairs within the bottleneck. Although one could argue that our methods lean towards a weak unlearning strategy–given the pre-trained backbone might retain some information about the forget set–our approach deviate from the strict definition of weak unlearning as outlined by Xu et al. (2023). As an example, when considering a non-parametric decoder, our methods affect intermediate rather than final model activations.
@@ 3 Background and Notations
Unlearning: Let $\mathcal{D}_{train}=\{x_{i},y_{i}\}^{N}_{i=1}$ be a training dataset and $\mathcal{D}_{test}$ be the corresponding test dataset. In our experiments, we consider the setting of class unlearning, wherein we aim to unlearn a class $c$ from a model trained with a multiclass classification objective on $\mathcal{D}_{train}$. $c$ is called the forget class or the forget set. Given $c$, we obtain $\mathcal{D}_{train}^{forget}\subset\mathcal{D}_{train}$ such that $\mathcal{D}_{train}^{forget}=\{(x,y)\in\mathcal{D}_{train}|y=c\}$. The complement of $\mathcal{D}_{train}^{forget}$ is $\mathcal{D}_{train}^{retain}$, i.e., subset of $\mathcal{D}_{train}$ that we wish to retain. Thus $\mathcal{D}_{train}^{retain}\cup\mathcal{D}_{train}^{forget}=\mathcal{D}_{train}$. Similarly, from $\mathcal{D}_{test}$, we have $\mathcal{D}_{test}^{forget}=\{(x,y)\in\mathcal{D}_{test}|y=c\}$ and its complement $\mathcal{D}_{test}^{retain}$. We refer to $\mathcal{D}_{train}^{retain}$ and $\mathcal{D}_{test}^{retain}$ as the retain set training and test data; and $\mathcal{D}_{train}^{forget}$ and $\mathcal{D}_{test}^{forget}$ as the forget set training and test data, respectively.
Discrete Key-Value Bottleneck: A discrete key-value bottleneck (DKVB) (Träuble et al., 2023) consists of a discrete set of coupled key-value codes. The bottleneck contains $C$ codebooks with each codebook containing $M$ key-value pairs. Models with DKVB use a pre-trained and frozen encoder to encode the input into a continuous representation. This input representation is then projected into $C$ lower dimension heads and each head is quantized to the $top-k$ nearest keys in the corresponding codebook. The values corresponding to the selected keys are averaged, and used for the downstream task. The keys in the codebooks are frozen and initialized to cover the input data manifold whereas the values are learnable. The mapping between the keys and values is non-parametric and frozen. Thus, the gradient is not propagated between the values and keys during training of the model. Since the values are retrieved and updated sparsely, and all the components except the value codes and the decoder are frozen, DKVB stores information in the form of input-dependent, sparse and localized representations (i.e., the value codes). These inductive biases allow the framework to exhibit improved generalization under distribution shifts during training, as shown empirically as well as theoretically in Träuble et al. (2023). Figure 1 (Left) shows an overview of a model with a DKVB where $C=1$, $M=5$ and top-$k=1$.